OpenStudy (anonymous):

(x − 7)2. solve please

OpenStudy (anonymous):

Distributive: 2x - 14

OpenStudy (anonymous):

2x-14

Parth (parthkohli):

No, it's squared.

OpenStudy (anonymous):

Just distribute it.. \[a(b - c) = ab - ac\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x - 7)(x - 7) }\) \( \color{Black}{\Rightarrow x(x - 7) - 7(x - 7)}\) \( \color{Black}{\Rightarrow x^2 - 7x - 7x + 49}\) \( \color{Black}{\Rightarrow x^2 - 14x + 49}\)

mathslover (mathslover):

let me give the answer of both : 1) \(\huge{2(x-7)=2(x)-2(7)=2x-14}\)

Parth (parthkohli):

You don't ever write it as (x - 7)2 but 2(x - 7), so it is squared. Believe me, I've seen a lot of questions like these.

OpenStudy (anonymous):

If it is square, then use the formula: \[(a-b)^2 = a^2 + b^2 - 2ab\] Here a = x and b = 7 Put them and solve it further...